OpenStudy (anonymous):

i just wan the answer and i have another question also Solve the system, using substitution. One point for correct steps, one point for correct ordered pair, (a,b). a + b = 4 and 2a + 3b = 11

OpenStudy (anonymous):

can someone give me the answer

OpenStudy (the_fizicx99):

Sorry, its against the CoC to give out answers.

OpenStudy (anonymous):

the what

OpenStudy (the_fizicx99):

@beccaboo333 <-- she goes full ham on this stuff

OpenStudy (anonymous):

is she on

OpenStudy (beccaboo333):

The code of conduct states that direct answers are not allowed. But instead we can help guide you to the answer alone.

OpenStudy (anonymous):

You can use 2 Rules for this type of problem to isolate one variable: 1. Add/subtract the two equations; 2. Substitution (Solve one equation for x and substitute it for x in the second equation)

OpenStudy (anonymous):

and who is gonna stop me from giving the answer

OpenStudy (beccaboo333):

Moderators.

OpenStudy (beccaboo333):

Eventually your school will deny you access back in.

OpenStudy (anonymous):

my school

OpenStudy (beccaboo333):

Administration from schools are on here, they do internal searches. If you cheat, you will be failed. If you keep doing it... You will be kicked out.

OpenStudy (anonymous):

Here, the question asks that you use substitution

OpenStudy (the_fizicx99):

Given that, a + b = 4 and 2a + 3b = 11 if you solve for A, a = -b + 4 2(-b + 4) + 3b = 11 -2b + 8 + 3b = 11 b + 8 = 11 -8 -8 b = 3

OpenStudy (the_fizicx99):

Solve for A using b

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

To substitute, isolate a or b

OpenStudy (anonymous):

f.e. a=4-b

OpenStudy (the_fizicx99):

@Tajar let him continue.

OpenStudy (anonymous):

SORRY!